• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name gaskell

    Log entry time 08:38:30 on October 22, 2010

    Entry number 336142

    keyword=Status update and plan for the day

    
    Yesterday, we completed a "Delta scan" acceptance test (see halog 336085)
    and on owl shift a deadtime test plan (see halog 336118). It looks like
    the ARS readout is in good shape but work remains to be done on the
    trigger (halog 336099).
    
    Plan for today:
    
    1. Change from 5 to 4 pass. This started at 8:30. 
    2. PD has requested we stay in beam permit until the pass is complete.
    Once MCC is done, we will take a Controlled Access. During that time, we
    will:
      a) Rotate the HRS to its new setting
      b) Fill the Moller
      c) Check a few issues with DAQ
    3. We will transition to production running at 4 pass (kinematics TBD)
    after the access is over.
    
    


    A copy of this log entry has been emailed to: munoz@jlab.org,hyde@jlab.org,jroche@jlab.org,camsonne@jlab.org